# On the Hurwitz Stability of Hurwitz-Type Matrix Polynomials

### v1

@@ Abstract
Every matrix polynomial ${\bf f}_{n}$ can be written in the form ${\bf f}_{n}(z)={\bf h}_{n}(z^{2})+z\,{\bf g}_{n}(z^{2})$. The matrix polynomial ${\bf f}_{2m}$ is said to be of Hurwitz-type if the expression ${\bf g}_{2m}(z){\bf h}_{2m}^{-1}(z)$ admits a representation as a finite continued fraction with positive definite matrix coefficients. Similarly, the odd-degree matrix polynomial ${\bf f}_{2m+1}$ is of Hurwitz type if $\frac{1}{z}{\bf h}_{2m+1}(z){\bf g}_{2m+1}^{-1}(z)$ has the same property. Using orthogonal matrix polynomials on the interval $[0,+\infty)$ and their associated second-kind polynomials, we establish the coprimeness of the matrix polynomial pairs $({\bf h}_{2m},{\bf g}_{2m})$ and $({\bf h}_{2m+1},z{\bf g}_{2m+1})$. Assuming a commutativity-type condition on the coefficients of $f_{n}$, we derive an explicit form of the Bezoutian associated with Hurwitz-type matrix polynomials. Under the same assumption, we also show that Hurwitz-type matrix polynomials are indeed Hurwitz matrix polynomials. Finally, we propose an extension of the class of Hurwitz-type matrix polynomials.
@@
Keywords: Hurwitz matrix polynomial, orthogonal matrix polynomial, Stieltjes transform, Bezoutian
@@ 1 Introduction
A scalar polynomial with all its roots in the left half of the complex plane is called a Hurwitz polynomial [60, 41, 58, 19, 32, 40, 46, 10, 42, 43]. The Hurwitz stability of such polynomials is also referred to as Hurwitzness—a term we adopt in this work. Let $n$ and $q$ be natural numbers. In the present work, we consider the properties and Hurwitzness of a subset of Lambda matrices [55, page 95], [33, page 228], [20], defined by
@@
${\bf h}_{n}(z)$ :=$:=\left\{\begin{array}[]{ll}A_{0}z^{m}+A_{2}z^{m-1}+\ldots+A_{2m},&n=2m,\\ A_{1}z^{m}+A_{3}z^{m-1}+\ldots+A_{2m+1},&n=2m+1,\end{array}\right.$ (5)
${\bf h}_{n}(z)$ :=$:=\begin{cases}A_{0}z^{m}+A_{2}z^{m-1}+\ldots+A_{2m},&n=2m,\\ A_{1}z^{m}+A_{3}z^{m-1}+\ldots+A_{2m+1},&n=2m+1,\end{cases}$ (6)
@@
${\bf g}_{n}(z)$ :=$:=\left\{\begin{array}[]{ll}A_{1}z^{m-1}+A_{3}z^{m-2}+\ldots+A_{2m-1},&n=2m,\\ A_{0}z^{m}+A_{2}z^{m-1}+\ldots+A_{2m},&n=2m+1.\end{array}\right.$ (9)
@@
The $q\times q$ matrix polynomial ${\bf f}_{n}$ in (1) is called a Hurwitz-type matrix polynomial if there exist two sequences of positive definite matrices, $({\bf c}_{k})_{k=0}^{m}$ and $({\bf d}_{k})_{k=0}^{m-1}$ (resp. $({\bf c}_{k})_{k=0}^{m}$ and $({\bf d}_{k})_{k=0}^{m}$) such that, for $n=2m$,
@@
The definition of HTM polynomials in the form of a finite continued fraction is inspired by [34, Chapter XV, Theorem 15].
The rational matrix functions in (10) and (11) are special solutions to the truncated Stieltjes matrix moment (TSMM) problem; see proof of [8, Theorem 7.9]. This implies that $\frac{{\bf g}_{n}(z)}{{\bf h}_{n}(z)}$ and $\frac{{\bf h}_{n}(z)}{z{\bf g}_{n}(z)}$ are Stieltjes transforms of a certain positive measure $\sigma$ on $[0,+\infty)$.
$s_{j}=\int\limits_{[0,+\infty)}t^{j}\sigma(dt),\quad 0\leq j\leq n-1,\quad s_{n}=\int\limits_{[0,+\infty)}t^{n}\sigma(dt)+M,$ (12)
where $M$ is a $q\times q$ complex valued positive semi-definite matrix.
$s(z)=\int_{[0,+\infty)}\frac{\sigma(dt)}{t-z},$ (13)
In the present work, we use the following condition on the coefficients of the matrix polynomial ${\bf f}_{n}$.
Condition C. For $n=2m$ (respectively, $n=2m+1$), assume that
$A_{2j+1}^{*}A_{2k}=A_{2k}^{*}A_{2j+1},$ (14)
for all $j=0,1,\ldots,m-1$ and $k=0,1,2,\ldots,m$ (respectively, $j=0,1,\ldots,m$ and $k=0,1,2,\ldots,m$).
This condition is a consequence of [2, Remark 2.2], which outlines the conditions for the Bezoutian of a rational matrix function to admit two distinct matrix factorizations. See also Remark 2.18.
@@
    We prove the coprimeness of the pair of matrix polynomials $({\bf h}_{2m}(z),{\bf g}_{2m}(z))$ (resp. $({\bf h}_{2m+1}(z),z{\bf g}_{2m+1}(z))$); see Theorem 3.5.
@@
    Under Condition C (14), we compute explicitly the Bezoutian of the HTM polynomial (1); see Theorem 4.4 and Corollary 4.5.
  • c)
    Under the same condition as in part b), we verify the Hurwitzness of HTM polynomials; see Theorem 5.3.
  • d)
    For a given matrix polynomial $P_{n}$ of degree $n$, which is not a Hurwitz-type matrix (HTM) polynomial, we construct a polynomial $f_{2n}(z)=P_{n}(z^{2})+zQ_{n-1}(z^{2})$ of degree $2n$, with a suitably chosen $Q_{n-1}$, such that ${\bf f}_{2n}$ becomes an HTM polynomial. Then, by applying the result from [25], we conclude that $P_{n}$ itself is an HTM polynomial; see Theorem 5.11.
The coprimeness of matrix polynomials was previously studied in [44, 66, 45, 3]. The Hurwitzness of matrix polynomials via Bezoutians was considered in [37, 53, 54]. Note that in [67] the Hurwitzness of the HTM polynomials was examined but without leveraging the commutative-type properties of the coefficients of the HTM polynomials. Moreover, the proof of [67, Theorem 3.5], which serves as the basis for the analysis of the Hurwitzness of HTM polynomials in [67], is not valid; see Remark 5.8.
Additionally, in contrast to [67], we use the representations of the matrix polynomial ${\bf f}_{n}$ as in (2), (74) and Condition C given by (14) to explicitly compute the Bezoutian of the quadruple (${\bf f}_{n}^{*}(\bar{x}),{\bf f}_{n}(-x),{\bf f}_{n}^{*}(-\bar{x}),{\bf f}_{n}(x)$). The asterisk denotes the Hermitian adjoint of a matrix, while the overbar represents complex conjugation. In the scalar case, the use of the ${\bf h}_{n}$ part and ${\bf g}_{n}$ part of the polynomial ${\bf f}_{n}$ for studying the corresponding Bezoutian was employed by Krein and Naimark in [48] and by Krein and Akhiezer in [1].
@@
Our work is organized as follows. In Section 2, we review previous results regarding HTM polynomials and their associated Markov parameters. Moreover, we mention Stieltjes nonnegative definite extendable sequences, Stieltjes orthogonal matrix polynomials relatively prime polynomial matrices, the Bezoutian, and the inertia of matrix polynomials. In Section 3, we examine the properties of HTM polynomials by exploring the relationships between the Markov parameters and the coefficients of HTM polynomials, as well as the interrelations between HTM polynomials and Markov parameters. Furthermore, we analyze the coprimeness of the matrix polynomial pairs $({\bf h}_{2m},{\bf g}_{2m})$ and $({\bf h}_{2m+1},z{\bf g}_{2m+1})$. In Section 4, we explicitly compute the Bezoutian of HTM polynomials. In Section 5, we study the Hurwitzness of HTM polynomials and we propose a procedure to derive an HTM from a given matrix polynomial.
@@ 2.1 HTM polynomials and associated Markov parameters
2.2 The Hankel nonnegative definite extendable sequences
Consider a TSMM problem with an odd (resp. even) number of moments $(s_{j})_{j=0}^{2m}$ (resp. $(s_{j})_{j=0}^{2m+1}$) as in (18). In [21], it is proven that the TSMM problem in the case of an odd (resp. even) number of moments has a solution if and only if $H_{1,m}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m}$) are both positive semi-definite matrices.
@@ 2.2 The Hankel nonnegative definite extendable sequences
We recall the concept of Stieltjes nonnegative definite extendable sequences for the TSMM problem for $n=2m+1$ and $n=2m$; refer to [30] for $\alpha=0$. See also [6], [27] and references therein.
Definition 2.3.
Let $(s_{j})_{j=0}^{2m-1}$ (resp. $(s_{j})_{j=0}^{2m}$) be a sequence from ${\mathbb{C}}^{q\times q}$ for which there exists some $s_{2m}\in{\mathbb{C}}^{q\times q}$ (resp. $s_{2m+1}\in{\mathbb{C}}^{q\times q}$) such that $H_{1,m}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m}$) are Hankel nonnegative definite block matrices. The sequence $(s_{j})_{j=0}^{2m-1}$ (resp. $(s_{j})_{j=0}^{2m}$) is called a Stieltjes nonnegative definite extendable sequence.
Next we paraphrase [30, Proposition 2.20] regarding Stieltjes nonnegative definite extendable sequences for $\alpha=0$.
Remark 2.4.
The set of Stieltjes positive sequences $(s_{j})_{j=0}^{2m}$ (resp. $(s_{j})_{j=0}^{2m+1}$) is a subset of the set of Stieltjes nonnegative definite extendable sequences $(s_{j})_{j=0}^{2m}$ (resp. $(s_{j})_{j=0}^{2m+1}$).
Let
$Y_{1,j}$ $:=y_{[j,2j-1]},\,\ 1\leq j\leq n,\,\quad\mbox{and}\quad Y_{2,j}:=y_{[j+1,2j]},\,\ 2\leq j\leq n,$ (19)
with
$y_{[j,k]}:=\left(s_{j},s_{j+1},\ldots,s_{k}\right)^{*},\quad 0\leq j\leq k$ (20)
and $y_{[j,k]}=0_{q}$ if $j>k$.
In the following proposition, we restate Propositions 4.13 and 4.14 from [30] for $\alpha=0$ using our notation. If $A$ and $B$ are $q\times q$ matrices, the inequality $A>B$ means that the matrix $A-B$ is positive definite.
Proposition 2.5.
a) The sequence $(s_{j})_{j=0}^{2m}$ is a positive Stieltjes sequence if and only if $(s_{j})_{j=0}^{2m-1}$ is a positive Stieltjes sequence and
$s_{2m}>Y_{1,m}^{*}H_{1,m-1}^{-1}Y_{1,m}.$ (21)
b) The sequence $(s_{j})_{j=0}^{2m+1}$ is a positive Stieltjes sequence if and only if $(s_{j})_{j=0}^{2m}$ is a positive Stieltjes sequence and
$s_{2m+1}>Y_{2,m}^{*}H_{2,m-1}^{-1}Y_{2,m}.$ (22)
Note that Proposition 2.5 indicates out the condition that matrices $s_{2m}$ and $s_{2m+1}$ must satisfy to be the extension of sequences $(s_{j})_{j=0}^{2m-1}$ and $(s_{j})_{j=0}^{2m}$, respectively.
@@ 2.3 Stieltjes orthogonal matrix polynomials
$T_{0}:=0_{q},\ \ T_{j}:=\left(\begin{array}[]{cc}0_{q\times jq}&0_{q}\\ I_{jq}&0_{jq\times q}\\ \end{array}\right),\qquad j\geq 1.$ (24)
@@
Furthermore, let $y_{[j,k]}$ be as in (20), and define
@@
Let $\widehat{H}_{1,j}$ (resp. $\widehat{H}_{2,j}$) denote the Schur complement of the block Hankel matrices $H_{1,j-1}$ in $H_{1,j}$ (resp.  $H_{2,j-1}$ in $H_{2,j}$) as seen in [57], [7, Equalities (2.5) and (2.6)]:
$\widehat{H}_{1,0}:=s_{0},\ \ \widehat{H}_{1,j}:=$ $s_{2j}-Y_{1,j}^{*}H_{1,j-1}^{-1}Y_{1,j},\ \ j\geq 1,$ (33)
$\widehat{H}_{2,0}:=s_{1},\ \ \widehat{H}_{2,j}:=$ $s_{2j+1}-Y_{2,j}^{*}H_{2,j-1}^{-1}Y_{2,j},\ \ j\geq 1.$ (34)
@@
The positive definite matrices in (35) to (37) of the following definition were introduced in [23] by Yu. Dyukarev.
@@
Let $(s_{j})_{j=0}^{2n}$ (resp. $(s_{j})_{j=0}^{2n+1}$) be a Stieltjes positive sequence. Furthermore, let $v_{j}$ and $u_{2,j}$ be as in (28) and (32), respectively. The $q\times q$ matrices
$M_{0}:=$ $s_{0}^{-1},\quad L_{0}:=s_{0}s_{1}^{-1}s_{0},$ (35)
$M_{j}:=$ $v_{j}^{*}H_{1,j}^{-1}v_{j}-v_{j-1}^{*}H_{1,j-1}^{-1}v_{j-1},$ (36)
$L_{j}:=$ $u_{2,j}^{*}H_{2,j}^{-1}u_{2,j}-u_{2,j-1}^{*}H_{2,j-1}^{-1}u_{2,j-1}$ (37)
are called the Dyukarev-Stieltjes matrix parameters of the TSMM problem.
T.J. Stieltjes’s renowned 1894 work “Recherches sur les fractions continues” [64] introduced the scalar version of the Dyukarev-Stieltjes parameters. These parameters, known as the Stieltjes parameters ([35], [49]), are given by $l_{j}:=\frac{\Delta_{j}^{2}}{\Delta_{j}^{(1)}\Delta_{j-1}^{(1)}}$, $m_{j}:=\frac{[\Delta_{j-1}^{(1)}]^{2}}{\Delta_{j}\Delta_{j-1}}$ and $j\geq 0$ with $\Delta_{j}:=\det(s_{i+k})_{i,k=0}^{j}$, $\Delta_{j}^{(1)}:=\det(s_{i+k+1})_{i,k=0}^{j}$, ($\Delta_{-1}:=\Delta_{-1}^{(1)}=1$).
Now we partially reproduce left orthogonal matrix polynomials (OMP) on $[0,+\infty)$ [25].
Definition 2.7.
Let $\sigma$ be a $q\times q$ nonnegenative measure on $[0,+\infty)$. A sequence $(P_{j})_{j=0}^{\infty}$ of $q\times q$ matrix-valued polynomials is called orthogonal with respect to $\sigma$ if
@@
$\int\limits_{[0,+\infty)}P_{j}(t)\,\sigma(dt)\,P_{k}^{*}(t)^{*}=\delta_{jk}C_{qj},$
@@
The matrix polynomials $P_{1,j}$ and $P_{2,j}$ are orthogonal with respect to a positive measure on $[0,+\infty)$. Simultaneously, the second-kind polynomials $Q_{1,j}$ and $Q_{2,j}$ can be derived from the corresponding polynomials $P_{1,j}$ and $P_{2,j}$, respectively.
@@
a) The polynomials $P_{1,j}$ and $P_{2,j}$ are OMP with respect to $\sigma$ and $t\sigma$, respectively. More precisely (as seen in [17, Remark 6.3]),
$\int\limits_{[0,+\infty)}P_{k,j}(t)t^{k-1}\sigma(dt)P_{k,l}^{*}(t)=\left\{\begin{array}[]{ccc}0_{q}&:&j\neq l\\ \widehat{H}_{k,j}&:&j=l\end{array}\right.,\ \ k=1,2.$ (43)
b) The following identities hold (as seen in [7, Remark D.6] and [7, Remark E.4]):
@@
$Q_{2,j}(x)=$ $\int\limits_{[0,+\infty)}\frac{xP_{2,j}(x)-tP_{2,j}(t)}{x-t}\sigma(dt),\ \ 0\leq j\leq n-1.$ (45)
By using the polynomials $P_{k,j}$ and $Q_{k,j}$ for $k=1,2$, we define the so-called resolvent matrix of the TSMM problem.
Definition 2.10.
[7, Theorem 4.7] For $k=0,1$, let $P_{k,j}$ and $Q_{k,j}$ be defined as in Definition 2.8. Define the $2q\times 2q$ matrix-valued function
$U^{(2m+k)}(z):=\left(\begin{array}[]{cc}Q_{2,m}^{*}(\bar{z})Q_{2,m}^{*^{-1}}(0)&-Q_{1,m+k}^{*}(\bar{z})P_{1,m+k}^{*^{-1}}(0)\\ -zP_{2,m}^{*}(\bar{z})Q_{2,m}^{*^{-1}}(0)&P_{1,m+k}^{*}(\bar{z})P_{1,m+k}^{*^{-1}}(0)\\ \end{array}\right),\quad k=0,1.$ (46)
For $k=0$ (resp. $k=1$), the matrix (46) is called the resolvent matrix of the TSMM problem for an odd number of moments $\left(s_{j}\right)_{j=0}^{2m}$ (resp. an even number of moments $\left(s_{j}\right)_{j=0}^{2m+1}$). In (46), the superscript of $U^{(2m+k)}$ for $k=0,1$ indicates the subscript of the moment $s_{2m+k}$ that appears in (46).
In [23], the identity
$U^{(2j)}(z)=$ $\left(\begin{array}[]{cc}I_{q}&0_{q}\\ -zM_{0}&I_{q}\end{array}\right)\left(\begin{array}[]{cc}I_{q}&L_{0}\\ 0&I_{q}\\ \end{array}\right)\ldots\left(\begin{array}[]{cc}I_{q}&L_{j-1}\\ 0_{q}&I_{q}\\ \end{array}\right)\left(\begin{array}[]{cc}I_{q}&0_{q}\\ -zM_{j}&I_{q}\end{array}\right)$ (55)
was verified and was used to express the solution of the TSMM problem through a finite matrix continued fraction. See the right-hand side of (60) and (61).
The complete set of solutions of the TSMM problem in the nondegenerate case is provided by the linear fractional transformation [17, Theorem 4.3]
$s^{(2m+k)}_{\scriptscriptstyle{\left(\begin{smallmatrix}{\bf p}\\ {\bf q}\end{smallmatrix}\right)}}(z):=\frac{Q_{2,m}^{*}(\bar{z})Q_{2,m}^{*^{-1}}(0)\,{{\bf p}}(z)-Q_{1,m+k}^{*}(\bar{z})P_{1,m+k}^{*^{-1}}(0)\,{{\bf q}}(z)}{-zP_{2,m}^{*}(\bar{z})Q_{2,m}^{*^{-1}}(0)\,{{\bf p}}(z)+P_{1,m+k}^{*}(\bar{z})P_{1,m+k}^{*^{-1}}(0)\,{{\bf q}}(z)},\,k=0,1,$ (56)
where ${\bf p}$ and ${\bf q}$ are certain meromorphic matrix-valued $q\times q$ functions; see [17, page 12]. The superscript of $s^{(2n+k)}_{\scriptscriptstyle{\left(\begin{smallmatrix}{\bf p}\\ {\bf q}\end{smallmatrix}\right)}}$ for $k=0,1$ indicates the subscript of the moment $s_{2m+k}$ appearing in (46), and the subscript of (56) emphasizes the dependency on the pair $\scriptscriptstyle{\left(\begin{smallmatrix}{\bf p}\\ {\bf q}\end{smallmatrix}\right)}$ for the associated solution $s^{(2m+k)}_{\scriptscriptstyle{\left(\begin{smallmatrix}{\bf p}\\ {\bf q}\end{smallmatrix}\right)}}$ to the TSMM problem.
By choosing the pair $\left(\begin{smallmatrix}{\bf p}\\ {\bf q}\end{smallmatrix}\right)$ equal to $\left(\begin{smallmatrix}I\\ 0\end{smallmatrix}\right)$ or $\left(\begin{smallmatrix}0\\ I\end{smallmatrix}\right)$, we attain the so-called extremal solutions to the TSMM problem. In particular, we will consider the following extremal solutions of the TSMM problem:
$s^{(2m)}_{\scriptscriptstyle{\left(\begin{smallmatrix}I\\ 0\end{smallmatrix}\right)}}(z):=-\frac{Q_{2,m}^{*}(\bar{z})}{z\,P_{2,m}^{*}(\bar{z})}$ (57)

### v3

@@ Abstract
Every matrix polynomial ${\bf f}_{n}$ can be written in the form
${\bf f}_{n}(z)={\bf h}_{n}(z^{2})+z\,{\bf g}_{n}(z^{2}).$
The matrix polynomial ${\bf f}_{2m}$ is said to be of Hurwitz type if the expression ${\bf g}_{2m}(z){\bf h}_{2m}^{-1}(z)$ admits a representation as a finite continued fraction with positive definite matrix coefficients. Similarly, the odd-degree matrix polynomial ${\bf f}_{2m+1}$ is of Hurwitz type if $\frac{1}{z}{\bf h}_{2m+1}(z){\bf g}_{2m+1}^{-1}(z)$ has the same property.
We derive an explicit form of the Bezoutian associated with Hurwitz-type matrix polynomials. Using this explicit form, we provide an explicit proof that Hurwitz-type matrix polynomials are Hurwitz matrix polynomials. In [52], the Hurwitzness of Hurwitz-type matrix polynomials was also studied.
Finally, we propose an extension of the class of Hurwitz-type matrix polynomials by adding to a non–Hurwitz-type matrix polynomial another matrix polynomial so that the resulting matrix polynomial is of Hurwitz type.
@@
Keywords: Hurwitz matrix polynomial, orthogonal matrix polynomial, Stieltjes transform, Bezoutian, Hurwitz-type matrix polynomial
@@ 1 Introduction
A scalar polynomial with all its roots in the left half of the complex plane is called a Hurwitz polynomial [50, 35, 48, 16, 28, 34, 38, 7, 36, 37]. The Hurwitz stability of such polynomials is also referred to as Hurwitzness—a term we adopt in this work. Let $n$ and $q$ be natural numbers. In the present work, we consider the properties and Hurwitzness of a subset of Lambda matrices [45, page 95], [29, page 228], [17], defined by
@@
${\bf h}_{n}(z)$ $:=\left\{\begin{array}[]{ll}A_{0}z^{m}+A_{2}z^{m-1}+\ldots+A_{2m},&n=2m,\\ A_{1}z^{m}+A_{3}z^{m-1}+\ldots+A_{2m+1},&n=2m+1,\end{array}\right$. (5)
@@
${\bf g}_{n}(z)$ $:=\left\{\begin{array}[]{ll}A_{1}z^{m-1}+A_{3}z^{m-2}+\ldots+A_{2m-1},&n=2m,\\ A_{0}z^{m}+A_{2}z^{m-1}+\ldots+A_{2m},&n=2m+1.\end{array}\right$. (8)
@@
The $q\times q$ matrix polynomial ${\bf f}_{n}$ in (1) is called a Hurwitz-type matrix polynomial if there exist two sequences of positive definite matrices, $({\bf c}_{k})_{k=0}^{m-1}$ and $({\bf d}_{k})_{k=0}^{m-1}$ (resp. $({\bf c}_{k})_{k=0}^{m-1}$ and $({\bf d}_{k})_{k=0}^{m}$) such that, for $n=2m$,
@@
The definition of HTM polynomials in the form of a finite continued fraction is inspired by [30, Chapter XV, Theorem 15].
@@
    We compute explicitly the Bezoutian of the HTM polynomial (1); see Theorem 3.4 and Corollary 3.5. From this, an explicit proof of the Hurwitz property of HTM polynomials readily follows; see Theorem 4.3.
@@
    For a given matrix polynomial $P_{n}$ of degree $n$ that is not a Hurwitz-type matrix (HTM) polynomial, we construct a polynomial ${\bf f}_{2n}(z)=P_{n}(z^{2})+zQ_{n-1}(z^{2})$ of degree $2n$, with a suitably chosen $Q_{n-1}$, such that ${\bf f}_{2n}$ becomes an HTM polynomial. Then, by applying the result from [22], we conclude that $P_{n}$ itself is an HTM polynomial; see Theorem 5.1.
The Hurwitz property of matrix polynomials via Bezoutians was studied in [31, 43, 44]. Note that in [52] the Hurwitz property of HTM polynomials was examined, but the proof of [52, Theorem 3.5], which serves as the basis for the analysis of the Hurwitz property of HTM polynomials in [52], is not fully detailed. More precisely, the proof of [52, Theorem 3.5] leaves certain steps implicit; see Remark 6.2. In particular, the argument is developed only for the case $n=2m$, whereas the case $n=2m+1$ is referred to as “similar,” without explicit justification.
Additionally, in contrast to [52], we use representations of the matrix polynomial ${\bf f}_{n}$ in terms of orthogonal polynomials on $0,\infty)$ of the first kind and polynomials of the second kind, as in (2) and (24), for example, to compute explicitly the Bezoutian of the quadruple $({\bf f}_{n}^{*}(\bar{x}),{\bf f}_{n}(-x),{\bf f}_{n}^{*}(-\bar{x}),{\bf f}_{n}(x))$.
The asterisk denotes the Hermitian adjoint of a matrix, while the overbar represents complex conjugation. In the scalar case, the use of the ${\bf h}_{n}$ part and ${\bf g}_{n}$ part of the polynomial ${\bf f}_{n}$ for studying the corresponding Bezoutian was employed by Krein and Naimark in [40] and by Krein and Akhiezer in [1].
@@
Our work is organized as follows. In Section 2, we review previous results regarding HTM polynomials and their associated Markov parameters. Moreover, we mention Stieltjes orthogonal matrix polynomials, orthogonal matrix polynomials, interrelations between HTM polynomials and Markov parameters, the Bezoutian, and the inertia of matrix polynomials.
In Section 3, we explicitly compute the Bezoutian of HTM polynomials. In Section 4, we study the Hurwitzness of HTM polynomials. In Section 5, we present a result and an algorithm for constructing an HTM polynomial from a given matrix polynomial that is not an HTM polynomial. Additionally, we indicate why the original matrix polynomial is a Hurwitz matrix polynomial. In Section 6, we explain in more detail certain steps left implicit in Theorem 3.5 of [52] on a “matrix analogue of the stability criterion via Markov parameters,” as stated on page 10 of [52].
@@ 2.1 HTM polynomials and associated Markov parameters
The rational matrix functions in (9) and (10) are special solutions to the truncated Stieltjes matrix moment (TSMM) problem; see proof of [5, Theorem 7.9]. This implies that $\frac{{\bf g}_{n}(z)}{{\bf h}_{n}(z)}$ and $\frac{{\bf h}_{n}(z)}{z{\bf g}_{n}(z)}$ are Stieltjes transforms of a certain positive measure $\sigma$ on $0,+\infty)$.
$s_{j}=\int\limits_{[0,+\infty)}t^{j}\sigma(dt),\quad 0\leq j\leq n-1,\quad s_{n}=\int\limits_{[0,+\infty)}t^{n}\sigma(dt)+M,$
where $M$ is a $q\times q$ complex valued positive semi-definite matrix.
$s(z)=\int_{[0,+\infty)}\frac{\sigma(dt)}{t-z},$
In [18], it is proven that the TSMM problem in the case of an odd (resp. even) number of moments has a solution if and only if $H_{1,m}$ and $H_{2,m-1}$ (resp. $H_{1,m}$ and $H_{2,m}$) are both positive semi-definite matrices.
@@ 2.2 Stieltjes orthogonal matrix polynomials
$T_{0}:=0_{q},\ \ T_{j}:=\left(\begin{array}[]{cc}0_{q\times jq}&0_{q}\\ I_{jq}&0_{jq\times q}\\ \end{array}\right),\qquad j\geq 1.$
@@
Furthermore, let
$Y_{1,j}:=y_{[j,2j-1]},\,\ 1\leq j\leq n,\,\quad\mbox{and}\quad Y_{2,j}:=y_{[j+1,2j]},\,\ 2\leq j\leq n,$
with
$y_{[j,k]}:=\left(s_{j},s_{j+1},\ldots,s_{k}\right)^{*},\quad 0\leq j\leq k$
and $y_{[j,k]}=0_{q}$ if $j>k$. Define
@@
Let $\widehat{H}_{1,j}$ (resp. $\widehat{H}_{2,j}$) denote the Schur complement of the block Hankel matrices $H_{1,j-1}$ in $H_{1,j}$ (resp.  $H_{2,j-1}$ in $H_{2,j}$) as seen in [47], [4, Equalities (2.5) and (2.6)]:
$\widehat{H}_{1,0}:=s_{0},\quad\widehat{H}_{1,j}:=$ $s_{2j}-Y_{1,j}^{*}H_{1,j-1}^{-1}Y_{1,j},\ \ j\geq 1,$
$\widehat{H}_{2,0}:=s_{1},\quad\widehat{H}_{2,j}:=$ $s_{2j+1}-Y_{2,j}^{*}H_{2,j-1}^{-1}Y_{2,j},\ \ j\geq 1.$
@@
Now we partially reproduce left orthogonal matrix polynomials (OMP) on $0,+\infty) [$22].
@@
Let $\sigma$ be a $q\times q$ nonnegenative measure on $0,+\infty)$. A sequence $(P_{j})_{j=0}^{\infty}$ of $q\times q$ matrix-valued polynomials is called orthogonal with respect to $\sigma$ if
@@
$\int\limits_{[0,+\infty)}P_{j}(t)\,\sigma(dt)\,P_{k}^{*}(t)=\delta_{jk}C_{qj},$
@@
In the next proposition, we recall that the matrix polynomials $P_{1,j}$ and $P_{2,j}$ are orthogonal with respect to a positive measure supported on $0,+\infty)$, and that the associated second-kind polynomials $Q_{1,j}$ and $Q_{2,j}$ can be obtained from $P_{1,j}$ and $P_{2,j}$, respectively. We also record an identity involving $P_{k,j}$ and $Q_{k,j}$ whose value is the zero matrix.
@@
a) The polynomials $P_{1,j}$ and $P_{2,j}$ are OMP with respect to $\sigma$ and $t\sigma$, respectively. More precisely (as seen in [14, Remark 6.3]),
$\int\limits_{[0,+\infty)}P_{k,j}(t)t^{k-1}\sigma(dt)P_{k,l}^{*}(t)=\left\{\begin{array}[]{ccc}0_{q}&:&j\neq l\\ \widehat{H}_{k,j}&:&j=l\end{array}\right.,\ \ k=1,2.$
b) The following identities hold (as seen in [4, Remark D.6] and [4, Remark E.4]):
@@
$Q_{2,j}(x)=$ $\int\limits_{[0,+\infty)}\frac{xP_{2,j}(x)-tP_{2,j}(t)}{x-t}\sigma(dt),\ \ 0\leq j\leq n-1.$
c) For $k=1,2$, the following identity holds (as seen in [4, Proposition 5.1 (a)])
$P_{k,j}(z)Q_{k,j}^{*}(\bar{z})-Q_{k,j}(z)P_{k,j}^{*}(\bar{z})=$ $0_{q}.$ (23)
For $n\geq 1$, it follows from [5, Theorem 6.1] that every HTM polynomial admits the representation
Identity (24) is obtained by using the relations